If the lateral area of a right circular cylinder is 90 cm2 and its volume is 225 cm3, find its radius.
1
Expert's answer
2020-12-08T07:47:22-0500

Area of right circular cylinder is "A= 2\pi r h" and volume is "V= \pi r^2 h" . Thus we have

"2\pi rh= 90 \\
rh =\cfrac{45}{\pi}......Eq[1]\\ 
r^2 h=\cfrac{225}{\pi}.......Eq[2]"

Now dividing Eq[2] by Eq[1]...

"r=\cfrac{225}{45}\\
r= 5cm.....Ans"

radius of right circular cylinder is 5 cm.